If you could count $1.46 per second, how many years would it take to count 1.00 billion dollars? eBook
Answer #1

Total amount = 1 billion dollars = 1*10^9 dollars

Counting speed = $1.46 per second

Number of seconds in a year = 3.154*10^7 sec

Amount counted in one year = Counting speed*number of seconds in a year

Amount counted in one year = 1.46*3.154*10^7 = 46048400 dollars

Now time taken to count 1 billion dollars in year

time taken = total amount/Amount counted per year

time taken = 1*10^9/46048400 = 21.7 years

Please Upvote.